OpenStudy (anonymous):

PRECAL! Is this correct about tan(π/16)? I know it shouldn't look exactly like this in one straight line...

OpenStudy (anonymous):

I know its really just a lot of plugging in... \[\tan \frac{ π }{ 16 }= \frac{ \sqrt{(1-\cosπ/8)} }{ \sqrt{(1+\cosπ/8)} } \]

OpenStudy (amistre64):

tan = sin/cos

OpenStudy (amistre64):

im not aware of any nice ways to address pi/16 tho.....

OpenStudy (anonymous):

\[\tan \frac{ π }{ 16 } = \frac{ \sin \frac{ π }{ 8 }}{ 1+\cos \frac{ π }{ 8 } }\]

OpenStudy (anonymous):

Is this the approach you were thinking of?

OpenStudy (amistre64):

\[tan(x)=\frac{sin(x)}{cos(x)}\] was the approach i was thinking of; but then there are half anlge and double anlge formulas that compound the issue

OpenStudy (anonymous):

Says to "use the half-angle formulas to come up with an exact expression..."

OpenStudy (amistre64):

the nice angles tend to be those constructed on the unit circle: 0,pi/3, pi/4, pi/6, and pi/2 trying to work these out into pi/16 can be a pain to say the least

OpenStudy (anonymous):

tan(π/16) = sin(π/8)/(1+cos(π/8)--------------------(1) Similarly, applying same formula to (π/4) tan(π/8) = sin(π/4)/(1+cos(π/4) = 1/(√2 +1) = √2 -1 sec² (π/8) = 2 +1-2√2 +1 = 2*(2-√2) or cos(π/8) = 1/√[2*(2-√2)] =[(1/4)*√{2*(2+√2)}]---------(2) and sin (π/8) =[{(√2 -1)/4}*√{2*(2+√2)}] ------------------(3) Substituting from (2) and (3) in (1), we get tan(π/16) =[{(√2 -1)/4}*√{2*(2+√2)}]/[1+[(1/4)*√{2*(2+√2)… tan(π/16) =[{(√2 -1)}*√{2*(2+√2)}]/[4+{√{2*(2+√2)}]

OpenStudy (amistre64):

i cant focus on these ....

OpenStudy (amistre64):

\[tan(\beta)=tan(\frac12\alpha)=\frac{tan(\alpha)}{1+sec(\alpha)}\] \[\frac{tan(\alpha)}{1+\sqrt{tan^2(\alpha)+1~}}\] \[tan(\alpha)=tan(\frac12\gamma)=\frac{tan(\gamma)}{1+sec(\gamma)}\] let \(\gamma\)=pi/4

OpenStudy (amistre64):

\[tan(\alpha) = \frac1{1+\sqrt2}\] \[tan^2(\alpha) = \frac1{3+2\sqrt2}\] \[tan(\beta)=\frac{1}{1+\sqrt{\cfrac1{3+2\sqrt2}+1~}}\] \[\frac{1}{1+\sqrt{\cfrac{1+3+2\sqrt2}{3+2\sqrt2}~}}\] \[\frac{1}{1+\sqrt{\cfrac{4+2\sqrt2}{3+2\sqrt2}~}}\] maybe ....

OpenStudy (anonymous):

okay... That was a different method